OpenStudy (anonymous):

Which of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) + 20? The graph of y = f(x) will shift left 20 units. The graph of y = f(x) will shift down 20 units. The graph of y = f(x) will shift up 20 units. The graph of y = f(x) will shift right 20 units.

OpenStudy (solomonzelman):

these are the rules of shifting the function \(\large\color{ blue }{\large {\bbox[5pt, lightgreen ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline~~~~~~~~~~~~~~~~~~~~~~~~~~~\textbf{Shifts}~~~~~~~~~~~~~~~~~~~~~~~~~~~&~\bf{c~~~units~~~~} \\ \hline \\f(x)= ∛x ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{red }{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= ∛x \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= ∛x \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

showing on a different example, but ...

OpenStudy (anonymous):

this chart is making no sense to me :S

OpenStudy (anonymous):

hold up so it would go up?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

20 units up

OpenStudy (anonymous):

cuz its addition with no square root

OpenStudy (solomonzelman):

yes, you are adding outside of the parenthesis (in other words)

OpenStudy (anonymous):

ur a genius :) thank you

OpenStudy (solomonzelman):

genius? not really, but you are very welcome:) Have fun !